OpenStudy (anonymous):

Candace bought 11 apples and 10 peaches at the grocery store. If the apples have a mass of 1,650 grams and the peaches have a mass of 1,200 grams, what is the total mass of the fruit? Need HelP!Thanks if you'll help me and I'll give you a medal.:)

OpenStudy (carniel):

Which fruit :P?

OpenStudy (anonymous):

a.2 kilograms 8 hectograms 5 dekagrams b.2 kilograms 8 centigrams 5 milligrams c.2 kilograms 8 grams 5 decigrams d.2 kilograms 8 dekagrams 5 grams

OpenStudy (carniel):

If you need to know just divide the numbers of fruit u have with the mass of it

OpenStudy (anonymous):

I think you need to convert all of that and the answers are in the choices.But I have no idea.:(

OpenStudy (whpalmer4):

Okay, if you add 1,650 grams and 1,200 grams, how many grams do you get?

OpenStudy (carniel):

Whats Kilo and Deca converted to as grams? K H D B d c m

OpenStudy (anonymous):

2,750 grams..right?

OpenStudy (anonymous):

2,850 grams i mean

OpenStudy (whpalmer4):

Right. Okay, how many grams in 1 kilogram?

OpenStudy (carniel):

>:O Whpalmer4!!

OpenStudy (carniel):

Seems I'm not needed ^^ GL

OpenStudy (anonymous):

1,000 g :)

OpenStudy (whpalmer4):

So if 1 kg is 1000 grams, how many kilograms can we take out of 2,850 grams, and what will we have left over?

OpenStudy (anonymous):

2.85

OpenStudy (whpalmer4):

How many whole kilograms?

OpenStudy (whpalmer4):

2 whole kilograms, right? That leaves 0.85 kilograms, or 850 grams. We need to figure out which combination of units will handle 850 grams.

OpenStudy (anonymous):

Oh okayy.:))

OpenStudy (whpalmer4):

This is where knowing the prefixes pays off. centi = 1/100 (1 cm = 1/100 m) hecto = 100 milli = 1/1000 (1 mm = 1/1000 m) deka = 10 deci = 1/10 what units will we need to make up 850 grams?